# Network-Realised Model Predictive Control Part II: Distributed Constraint Management

### v4

@@ 4.2 Nominal Area-based Prediction Models
In order to formulate tractable prediction models for our architecture’s second layer, we need to account for the state variables of the implementations from (10). Thus, aside from the $x$ and $u$ subscripts appearing in (5)-(6), we introduce the additional subscript $w$ to refer to the state vector of each area’s NRF subcontroller, and we extend the index collection defined originally in (4) to $\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui},\mathcal{A}_{wi}),\,\forall\,i\in\{1:N\}$.
@@
where we have that
@@
    the residual contributions of initial conditions from outside of the $i^{\text{th}}$ area’s neighbourhood and of cross-coupling with other areas’ MPC subcontrollers is
@@
The quantities described in (19) are called residual because it is the primary goal of the first layer to attenuate the effect that these signals have on the area dynamics from (16a)-(16d). Since these terms, along with those in (17), represent the contribution of variables which are not available (in terms of measurement or prediction) to the $i^{\text{th}}$ area, the only sensible recourse is to minimise their influence (see Section 5 in [part1]).
@@ 4.4 Connecting the Augmented Prediction Models to the Nominal Ones
Since the vector $\tiny\begin{bmatrix}\nu_{x}^{\top}&\nu_{w}^{\top}\end{bmatrix}^{\top}$ cannot be determined a priori, our MPC-based subcontrollers will have to make use of the initial conditions in (20) instead of the real ones. Consequently, our control policies will employ the prediction models affected by additive disturbance (21a)-(21b), instead of the ones given in (16b)-(16c). However, a straightforward computation yields the fact that the original area dynamics, given by the identities obtained in (16b)-(16c), can be retrieved from (21a)-(21b) via
$x_{i}[k]=\widetilde{x}_{i}[k]-\eta_{xi}[k],\quad u_{fi}[k]=\widetilde{u}_{fi}[k]-\eta_{ui}[k],\vskip-8.53581pt$ (24)
@@
Having obtained the area-based prediction models in (16a)-(16d) and having augmented them as in (21a)-(21b), we now proceed to discuss constraint formulation for the optimisation problems described in (8).
@@ 5 Constraints in the Distributed Setting
The assumption that the pair $(\mathcal{X},\mathcal{U})$ can be decomposed as in (26) is not only required, when attempting to break down the global problem into any configuration of local subproblems, but it is also a mild one (see its equivalent counterpart in [DMPC1, DMPC2]). Indeed, the conditions located on the left-hand side of (26) are often represented (in practice) by box constraints, for which said decomposition may always be performed.
@@ 5.1 Placing Constraints on Exogenous Disturbance
One of the key benefits of our two-layer approach is the fact that, from the point of view of the NRF-based layer, the signals $u_{s1}[k]$ and $u_{s2}[k]$ are perceived in the same manner as the all the other components the vector from (7), namely as additive and exogenous disturbance. Consequently, the special disturbance attenuation properties (such as string stability, see [plutonizare] and [SS_rev], along with the references therein) of our control system are preserved when adding the second layer on top of the first. Note that such special properties are particularly challenging to ensure using exclusively MPC-based techniques, a topic which has generated a great deal of research in system-theoretical literature [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5]. One of the main contributions of the present work is the fact that it combines the native benefits of two separate techniques, disturbance attenuation for NRF and constraint management for MPC, into a single framework.
@@ 5.3 Placing Constraints on First-Layer States
Since the initial conditions of the first layer’s state-space implementation explicitly impact the signals located on the left-hand side of (14), it follows that the state vectors of the NRF-based subcontrollers from (10) should also possess a state-based description, which will be used to ensure the satisfaction of the inclusions in (26).
@@
The results presented in this section represent one of the key theoretical novelties introduced in this paper. In the classical tube MPC setting (see, for example, [R8A, R8B, R8D]), the command constraints on the signals analogous to $u_{fi}[k]$ are straightforward to compute by means of constant matrices. In our framework, however, it becomes necessary to pass through the dynamics in (10), which significantly complicates the resulting expressions. Given the delicate nature of this undertaking, said expressions are presented, within the current section, in as explicit a manner as possible, regardless of the apparent similarity with standard tube MPC (recall the first paragraph of Section 4). Nevertheless, the benefits of NRF-based control (as stated in Remark 5.2) more than make up for the increase in complexity when formulating the synthesis procedure (and not, we point out, when implementing the control laws). Indeed, a further advantage of our approach is the fact that, instead of limiting ourselves to block-diagonal and constant stabilising feedback (see Assumption 1 in [R8A, R8B, R8D]), we leverage a greater degree of freedom in considering dynamical stabilising feedback (10) endowed with generic sparsity patterns (11).
@@
In order to state the main result of this section, we introduce the sets ${\mathcal{W}}_{\ell 1}\subseteq\mathbb{R}$, with $\ell\in\{1:n_{u}\}$, which satisfy
@@
for all $i\in\{1:N\}$. The following result employs the newly introduced sets with the aim of constraining the state variables of the first-layer subcontrollers.
@@
Let the first-layer subcontrollers be implemented as in (9)-(13). For each $i\in\{1:N\}$, define ${Z}_{ri}:=\mathrm{diag}(S_{ui},S_{xi})$ and the sets from (31) and (34a)-(34b), located at the top of this page. Assume that:
@@ 3. A3)
Note that the $i^{\text{th}}$ area’s sets given in (34a)-(34b) are computed efficiently, using only for those indices belonging to $\mathcal{N}_{i}$. This is owed to the sparsity structure displayed in (13), which serves to incentivise the synthesis of sparse control laws for the first layer (see Appendix A, along with [part1]). Note that the bounds associated with these sets could also be tightened by adapting the ideas presented in [R8B], through the sequential reduction of those $\mathcal{X}_{j}$, $\mathcal{U}_{j}$, $\mathcal{U}_{s1j}$ and $\mathcal{U}_{s2j}$ for which $j\in\mathcal{N}_{i}$.
@@ 5.4 Addressing Uncertainty in the Initial Conditions
Thus, our problem reduces to employing the dynamics from (21a)-(21b) with the aim of constraining $\widetilde{x}_{i}[k]$ and $\widetilde{u}_{fi}[k]$ to the sets obtained in (36) and (37), respectively. Recalling the identities from (24) directly yields the desired inclusions from (26). All that remains is to compute bounding sets for the signals from (22)-(23).
@@ 5.5 Placing Constraints on Area Cross-coupling
In order to provide compact formulations in the sequel, we first construct the sets $\mathcal{W}_{i}\subseteq\mathbb{R}^{n_{wi}}$ which satisfy
@@
for all $i\in\{1:N\}$, by simply taking the Cartesian product of the $\mathcal{W}_{\ell j}$ sets appearing in (38). Since the aim of the second layer is to constrain the state variables of the network and those of the implementations given in (10) to the sets $\mathcal{X}_{i}$ and $\mathcal{W}_{i}$, respectively, we consider that the initial conditions of the first layer’s closed-loop system satisfy $x_{ci}\in\mathcal{X}_{i}$ along with $w_{ci}\in\mathcal{W}_{i}$, for all $i\in\{1:N\}$. In this case, the definitions introduced in (20) yield
@@
Note that, in choosing the sets from (30b), one negotiates a command-budgeting balance (as limited by $\mathcal{U}_{i}$) between the first and second layers. This induces a trade-off associated with selecting the sets from (30a)-(30b), since the more freedom to act $u_{s1i}[k]$ is given, via the set from (30a), its effect will expand throughout the sets from (31) and (34a)-(34b). Similarly, if the network’s areas are not all perfectly decoupled (recall Remark 4.4), then too much leeway given to $u_{s1i}[k]$ and $u_{s2i}[k]$ may result in significant area cross-coupling. In this case, one may derive an extended version of the technique from [R8C] (which manifests its procedure’s effectiveness on only two collaborating areas) to mitigate these particular aspects. Section 4 in [R8C] also provides valuable insight into negotiating this particular type of trade-off.
@@ 6 Distributed Implementation of the MPC Layer
In this section, we proceed to show how to ensure global theoretical guarantees for the first layer’s closed-loop system, by independently implementing an MPC controller for each area’s dynamics described in (16a)-(16d). The results presented here are inspired by the core idea investigated in [AI], and our proposed technique represents the MPC-oriented analogue of the indicated work.
@@
As discussed in this manuscript’s companion paper [part1], our approach showcases a series of notable similarities with the SLS-based procedure from [DMPC1, DMPC2]. Indeed, in both instances, the focus rests upon designing the closed-loop response of the network in accordance with a number of structural restrictions. Additionally, the d-locality constraints from [DMPC1, DMPC2] closely match our own area-based sparsity patterns discussed in Remark 4.4, while the feasibilities of these two approaches are also closely intertwined, since both the SLS [SLS] and the Youla Parametrisation (upon which the first layer is based, see [NRF]) yield the set of all achievable and stable closed-loop maps (see [Luca3]). On the other hand, our procedure leverages the fact that the NRF-based layer can always be designed and implemented offline, thus relegating all online computation to the significantly less resource-demanding MPC layer, as shown in the sequel. This stands in stark contrast to the approach from [DMPC1, DMPC2], where sparsity constraints must be factored into the online computation, alongside all the MPC-like constraints.
@@ 6.1 Distributed Enforcement of Constraint Satisfaction
We begin by addressing the final and most important component of the formulations introduced in (8). Recall that, in stating the objectives of our control architecture’s second layer, the optimisation problems from (8) make use of the sets $\Xi_{it}$ to ensure the satisfaction of the constraints pertaining to $x[k]$ and $u[k]$. The aforementioned sets, whose precise expressions have been deferred up to this point, may now finally be stated in a formal manner. This is done by first computing the sets
@@
We implicitly assume that the first-layer closed-loop system has been designed in such a way as to ensure that the sets from (42)-(44) are non-empty. If this is not the case, then the reason is the fact that the first layer has not been appropriately tuned in order to tackle cross-coupling and the exogenous disturbance. Provided that the sets $\mathcal{X}$ and $\mathcal{U}$ (along with their area-based decompositions) are available during the first layer’s design phase, a judicious choice of systemic norm optimisation in Section 5.2 of [part1] (tied to the nature of these sets) can be used to ensure non-emptiness. Indeed, given that our first layer is specifically designed to improve upon the open-loop response characteristics of the network described by (1a)-(1c), our assumption is by no means restrictive. Moreover, it highlights one of the key benefits of our approach: the fact that constraint feasibility is certified after first-layer synthesis, with any potential infeasibility being detected early on during the design procedure.
With all of their components now clarified, and assuming that every problem described in (8) is feasible [^3] This property will be rigorously addressed in the sequel. for each time instant $k\geq k_{0}$, each MPC-based subcontroller in our second layer can be implemented as per Algorithm 1.
@@
It is straightforward to notice that the tuples formed by $(H_{1it},h_{1it})$ and $(H_{2\ell t},h_{2\ell t})$ characterise polyhedral inner-approximations of the sets on the right-hand sides of (42)-(43). Crucially, these pairs can always be determined offline, which greatly simplifies the online construction of the sets from (44) in Step 3 of Algorithm 1. In addition to this computational benefit, note that no explicit construction of any invariant sets occurs in our setting. For especially large or highly interconnected networks, the type of invariant sets discussed in [R8A] or [R8C] may be highly complex, or even numerically intractable for implementation. Our framework allows us to bypass explicit invariance-related concerns, and it allows us to control the complexity of the sets in (44), by means of the inner approximations given in (42)-(43).
@@
Regarding the computation of said command signals, the MPC formulations provided in (8) are particularly flexible, in the context of (distributed) constraint satisfaction. The two prediction horizons, introduced in point c) of Section 3.2 and associated with these problems, permit the careful management of the computational burdens which come with solving (8). Since we are chiefly interested in satisfying the latter’s constraints only for the sake of recursive feasibility (see the numerical example presented in the sequel), it is possible to improve closed-loop performance of the MPC-based layer by increasing the unconstrained prediction horizon without dramatically impacting the overall computational cost.
@@ 2. A2)
Note that Assumption A1) from Proposition 6.1 hinges upon the initial condition of the controlled network and upon the sets $\mathcal{X}_{i}$ in (26), since the initial conditions of the implementations in (10) can be imposed at will. Should the system from (1a)-(1c) be initialised in some $x_{nf}\not\in\mathcal{X}$, the procedure presented in this paper and in [part1] can also be used as a recovery mechanism. The key is to relax the state constraints given by $\mathcal{X}$ into some feasible and connected set $\mathcal{X}_{r}\supseteq(\mathcal{X}\cup\{x_{nf}\})$, while keeping the same $\mathcal{U}$ and resynthesising the second layer in accordance with the relaxed constraints. Since, in our framework, the first layer is the one responsible for disturbance rejection and reference tracking, the NRF-based control laws can then be used to drive the network’s state within $\mathcal{X}$, all while the second layer handles command limitation. The first layer’s subcontrollers can then be reinitialised to satisfy Assumption A1), and nominal operation may commence using the original second layer.
When starting from a point inside the state-constraining set $\mathcal{X}$, Proposition 6.1 shows that our control scheme is able to maintain the first layer’s state variables inside $\mathcal{X}$ while also satisfying the command constraints encoded by $\mathcal{U}$, provided that all of the MPC-based problems from (8) are feasible. Given the critical importance of satisfying Assumption A2) from Proposition 6.1, we now proceed to investigate a set of conditions which offer strong theoretical guarantees regarding the feasibility of the optimisation problems from (8).

### v5

@@ 4.2 Nominal Area-based Prediction Models
To formulate tractable prediction models for our architecture’s second layer, we need to account for the state variables of the implementations given in (10). Therefore, aside from the $x$ and $u$ subscripts appearing in (5)-(6), we introduce the additional subscript $w$ to refer to the state vector of each area’s NRF subcontroller, and we extend the index collection defined initially in (4) to
$\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui},\mathcal{A}_{wi}),\,\forall\,i\in\{1:N\}.\vskip-5.69054pt$
@@
where we have that:
@@
    the residual contribution of initial conditions from outside of the $i^{\text{th}}$ area’s neighbourhood and of cross-coupling with other areas’ MPC subcontrollers is
@@
The quantities described in (19) are called residual because it is the primary goal of the first layer to attenuate the effect that these signals have on the area dynamics from (16a)-(16d). Since these terms, along with those in (17), represent the contribution of variables which are not available (in terms of measurement or prediction) to the $i^{\text{th}}$ area, the only sensible recourse is to minimise their influence (see Section 5 in [24]).
@@ 4.4 Connecting the Augmented Prediction Models to the Nominal Ones
Since the vector $\tiny\begin{bmatrix}\nu_{x}^{\top}&\nu_{w}^{\top}\end{bmatrix}^{\top}$ cannot be determined a priori, our MPC-based subcontrollers will have to use the initial conditions in (20) instead of the actual ones. Consequently, our control policies will employ the prediction models affected by additive disturbance (21a)-(21b), instead of the ones given in (16b)-(16c). However, a straightforward computation yields the fact that the original area dynamics, given by the identities obtained in (16b)-(16c), can be retrieved from (21a)-(21b) via
$x_{i}[k]=\widetilde{x}_{i}[k]-\eta_{xi}[k],\quad u_{fi}[k]=\widetilde{u}_{fi}[k]-\eta_{ui}[k],\vskip-5.69054pt$ (24)
@@
Having obtained the area-based prediction models in (16a)-(16d) and having augmented them as shown in (21a)-(21b), we now proceed to discuss constraint formulation for the optimisation problems described in (8).
@@ 5 Constraints in the Distributed Setting
The assumption that the pair $(\mathcal{X},\mathcal{U})$ can be decomposed as in (26) is not only required, when attempting to break down the global problem into any configuration of local subproblems, but it is also a mild one (see its equivalent counterpart in [1, 2]). Indeed, the conditions located on the left-hand side of (26) are often represented (in practice) by box constraints, for which said decomposition may always be performed.
@@ 5.1 Placing Constraints on Exogenous Disturbance
One of the key benefits of our two-layer approach is the fact that, from the point of view of the NRF-based layer, the signals $u_{s1}[k]$ and $u_{s2}[k]$ are perceived in the same manner as the all the other components of the vector from (7), namely as additive and exogenous disturbance. Consequently, the special disturbance attenuation properties (such as string stability, see [20] and [9], along with the references therein) of our control system are preserved when adding the second layer on top of the first. Note that such special properties are particularly challenging to ensure using exclusively MPC-based techniques, a topic which has generated a great deal of research in system-theoretical literature [31, 28, 6, 10, 27]. One of the main contributions of the present work is that it combines the native benefits of two separate techniques, disturbance attenuation for NRF-based control and constraint management for MPC, into a single framework.
@@ 5.3 Placing Constraints on First-Layer States
Since the initial conditions of the first layer’s state-space implementation explicitly impact the signals located on the left-hand side of (14), it follows that the state vectors of the NRF-based subcontrollers from (10) should also possess a set-based description, which will be employed to ensure the satisfaction of the inclusions from (26).
@@
The results presented in this section represent one of the key theoretical novelties introduced in this paper. In the classical tube MPC setting (see, for example, [8, 26, 11]), the command constraints on the signals analogous to $u_{fi}[k]$ are straightforward to compute by means of constant matrices. In our framework, however, it becomes necessary to pass through the dynamics in (10), which significantly complicates the resulting expressions. Given the delicate nature of this undertaking, said expressions are presented, within the current section, in as explicit a manner as possible, regardless of the apparent similarity with standard tube MPC (recall the first paragraph of Section 4). Nevertheless, the benefits of NRF-based control (as stated in Remark 5.2) more than make up for the increase in complexity when formulating the synthesis procedure (and not, we point out, when implementing the control laws). Indeed, a further advantage of our approach is the fact that, instead of limiting ourselves to block-diagonal and constant stabilising feedback (see Assumption 1 in [8, 26, 11]), we leverage a greater degree of freedom in considering dynamical stabilising feedback (10) endowed with generic sparsity patterns (11).
@@
To state the main result of this section, we introduce the sets ${\mathcal{W}}_{\ell 1}\subseteq\mathbb{R}$, with $\ell\in\{1:n_{u}\}$, which satisfy
@@
for all $i\in\{1:N\}$. The following result employs the newly introduced sets to constrain the state variables of the first-layer subcontrollers.
@@
Let the first-layer subcontrollers be implemented as in (9)-(13). For each $i\in\{1:N\}$, define ${Z}_{ri}:=\mathrm{diag}(S_{ui},S_{xi})$ along with the sets from (28) and (34a)-(34b), which are located at the top of this page. Assume, moreover, that:
@@ 3. A3)
Note that the $i^{\text{th}}$ area’s sets given in (34a)-(34b) are computed efficiently, using only for those indices belonging to $\mathcal{N}_{i}$. This efficient computation is owed to the sparsity structure displayed in (13), which incentivises the synthesis of sparse control laws for the first layer (see Appendix A and [24]). Note that the bounds associated with these sets could also be tightened by adapting the ideas presented in [26], through the sequential reduction of those $\mathcal{X}_{j}$, $\mathcal{U}_{j}$, $\mathcal{U}_{s1j}$ and $\mathcal{U}_{s2j}$ for which $j\in\mathcal{N}_{i}$.
@@ 5.4 Addressing Uncertainty in the Initial Conditions
Thus, our problem reduces to employing the dynamics from (21a)-(21b) with the aim of constraining $\widetilde{x}_{i}[k]$ and $\widetilde{u}_{fi}[k]$ to the sets obtained in (36) and (37), respectively. Recalling the identities from (24) directly yields the desired inclusions stated in (26). All that remains is to compute containing sets for the signals from (22)-(23).
@@ 5.5 Placing Constraints on Area Cross-coupling
To provide compact formulations in the sequel, we first construct a collection of sets $\mathcal{W}_{i}\subseteq\mathbb{R}^{n_{wi}}$ which satisfy
@@
for all $i\in\{1:N\}$, by simply taking the Cartesian product of the $\mathcal{W}_{\ell j}$ sets appearing in (38). Since the chief goal of the second layer is to constrain the state variables of the network and those of the implementations given in (10) to the sets $\mathcal{X}_{i}$ and $\mathcal{W}_{i}$, respectively, we consider that the initial conditions of the first layer’s closed-loop system satisfy $x_{ci}\in\mathcal{X}_{i}$ along with $w_{ci}\in\mathcal{W}_{i}$, for all $i\in\{1:N\}$. In this case, the definitions from (20) yield
@@
In choosing the sets from (31b), one negotiates a command-budgeting balance (as limited by $\mathcal{U}_{i}$) between the first and second layers. This choice induces a trade-off associated with selecting the sets from both (31a) and(31b), since the more freedom to act $u_{s1i}[k]$ is given, via the set from (31a), its effect will expand throughout the sets from (28) and (34a)-(34b). Similarly, if the network’s areas are not all perfectly decoupled (recall Remark 4.4), then too much leeway given to $u_{s1i}[k]$ and to $u_{s2i}[k]$ may result in significant area cross-coupling. In this case, one may derive an extended version of the technique from [22] (which demonstrates its effectiveness for only two collaborating areas), so as to mitigate these particular aspects. Section 4 in [22] also provides valuable insight into negotiating this particular type of trade-off.
@@ 6 Distributed Implementation of the MPC Layer
In this section, we proceed to show how to ensure global theoretical guarantees for the first layer’s closed-loop system, by independently implementing an MPC controller for each area’s dynamics described in (16a)-(16d). The results presented here are inspired by the core idea investigated in [17], and our proposed technique represents the MPC-oriented analogue of the indicated work.
@@
As discussed in this manuscript’s companion paper [24], our approach exhibits several notable similarities with the SLS-based procedure from [1, 2]. Indeed, in both instances, the focus rests upon designing the network’s closed-loop response in accordance with a number of structural restrictions. Additionally, the d-locality constraints from [1, 2] closely match our own area-based sparsity patterns discussed in Remark 4.4, while the feasibilities of these two approaches are also closely intertwined, since both the SLS [29] and the Youla Parametrisation (upon which the first layer is based, see [21]) yield the set of all achievable and stable closed-loop maps (see also [30]). On the other hand, our procedure leverages the fact that the NRF-based layer can always be designed and implemented offline, thus relegating all online computation to the significantly less resource-demanding MPC layer, as shown in the sequel. This feature stands in stark contrast to the framework presented in [1, 2], in which sparsity constraints must be factored into the online computation, alongside all of the MPC-like constraints.
@@ 6.1 Distributed Enforcement of Constraint Satisfaction
We begin by addressing the final and most important component of the formulations introduced in (8). Recall that, in stating the objectives of our control architecture’s second layer, the optimisation problems from (8) make use of the sets $\Xi_{it}$ to ensure the satisfaction of the constraints related to $x[k]$ and $u[k]$. The aforementioned sets, whose precise expressions have been deferred up to this point, may now finally be stated in a formal manner. To do so, we first compute the following sets
@@
We implicitly assume that the first-layer closed-loop system has been designed so as to ensure that the sets from (42)-(44) are non-empty. If this is not the case, the reason is that the first layer has not been appropriately tuned to address cross-coupling and exogenous disturbance. Provided that the sets $\mathcal{X}$ and $\mathcal{U}$ (along with their area-based decompositions) are available during the first layer’s design phase, a judicious choice of systemic norm optimisation in Section 5.2 of [24] (tied to the nature of these sets) can be used to ensure non-emptiness. Indeed, given that our first layer is specifically designed to improve upon the open-loop response characteristics of the network described by (1a)-(1c), our assumption is by no means restrictive. This fact also highlights one of the key benefits of our approach: constraint feasibility is certified after first-layer synthesis, with any potential infeasibility being detected early on in the design procedure.
With all of their components clarified, and assuming that every problem described in (8) is feasible for each time instant $k\geq k_{0}$ (a key property that will be rigorously addressed in the sequel), each MPC-based subcontroller in our second layer can be implemented as per Algorithm 1.
@@
It is straightforward to notice that the tuples formed by $(H_{1it},h_{1it})$ and $(H_{2\ell t},h_{2\ell t})$ characterise polyhedral inner-approximations of the sets on the right-hand sides of (42)-(43). Crucially, these pairs can always be determined offline, which greatly simplifies the online construction of the sets from (44) in Step 3 of Algorithm 1. In addition to this computational benefit, note that no explicit construction of any invariant sets occurs in our setting. For especially large or highly interconnected networks, the type of invariant sets discussed in [8] or [22] may be highly complex, or even numerically intractable for practical implementations. Our framework allows us to bypass explicit invariance-related concerns, and it allows us to control the complexity of the sets in (44), by means of the inner approximations in (42)-(43).
@@
Regarding the computation of said command signals, the MPC formulations provided in (8) are particularly flexible, in the context of (distributed) constraint satisfaction. The two prediction horizons, introduced in point c) of Section 3.2 and associated with these problems, permit the careful management of the computational burdens which come with solving (8). Since we are chiefly interested in satisfying the latter’s constraints only for the sake of recursive feasibility (see the numerical example presented in the sequel), it is possible to improve the closed-loop performance of the MPC-based layer by increasing the unconstrained prediction horizon without significantly impacting the overall computational cost.
@@ 2. A2)
Note that Assumption A1) from Proposition 6.1 hinges upon the initial condition of the controlled network and upon the sets $\mathcal{X}_{i}$ in (26), since the initial conditions of the implementations in (10) can be imposed at will. Should the system from (1a)-(1c) be initialised in some $x_{nf}\not\in\mathcal{X}$, the procedure presented in this paper and in [24] can also be used as a form of recovery mechanism. The key is to relax the state constraints given by $\mathcal{X}$ into some feasible and connected set $\mathcal{X}_{r}\supseteq(\mathcal{X}\cup\{x_{nf}\})$, while keeping the same $\mathcal{U}$ and resynthesising the second layer in accordance with the relaxed constraints. Since, in our framework, the first layer is responsible for disturbance rejection and reference tracking, the NRF-based control laws can then be used to drive the network’s state within $\mathcal{X}$, while the second layer handles command limitation. The first layer’s subcontrollers can then be reinitialised so as to satisfy Assumption A1), and nominal operation may commence using the original second layer.
When starting from a point inside the state-constraining set $\mathcal{X}$, Proposition 6.1 shows that our control scheme is able to maintain the first layer’s state variables inside $\mathcal{X}$ while also satisfying the command constraints encoded by $\mathcal{U}$, provided that all of the MPC-based problems showcased in (8) are feasible. Thus, given the critical importance of satisfying Assumption A2) from Proposition 6.1, we now proceed to investigate a set of conditions which offer strong theoretical guarantees regarding the feasibility of the optimisation problems from (8).